Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, mwas, bersanetti - 22:33 Tuesday 18 November 2025 (68188) Print this report
ITF recovery

This afternoon/evening we worked on the recovery of the ITF.

Over the afternoon we observed multiple unlocks in different points of the lock acquisition that we ultimately discovered being caused by a mistuned phase for B2_169MHz. Looking at the MICH_PRCL_COUPLING channel, this phase was re-tuned along the lock acquisition and CARM_NULL_1F can now be reached smoothly.

Here we found new issues in locking the OMC: a bright fringe forces to misalign the SR (TY +2 urad), which in turn introduces an imbalance in the sidebands. Without a MICH offset, this imbalance makes B1p shoot up when the DARM offset is introduced and the ITF unlocks while attempting to lock the OMC.

Introducing both a misalignment of +2 urad on SR_TY and a MICH offset of 40 produces a condition in which the OMC can be locked.

With yet another stroke of bad luck, B1_PD1 is not working and does not respond to the rearm command, preventing us from reaching DC_READOUT.

This issue cannot be solved tonight, but at least we recovered the lock acquisition up to CARM_NULL_1F. We leave the ITF in this state.

Comments to this report:
mwas - 8:39 Wednesday 19 November 2025 (68189) Print this report

One more thing that happened yesterday afternoon is that the power on the B1s camera was several orders of magnitude higher than normal. This was likely a consequence of trying to recover the OMC slow shutter in the morning, one of the waveplates of the OMC must have been rotated. A check in single bounce has confirmed that most of the power on B1s is in the P instead of S polarization. Moving the north waveplate by ~10'000 steps (half-wave plate if I remember well) recovered the good polarization. I haven't done a fine tuning, and the south wave plate might be off by ~100 steps, as I had first done small steps with both waveplates to try to understand which one was moved by mistake this morning (sending two commands in a row too close together can confuse the driver and make it turn a different waveplate).

This morning Alain has recovered B1 PD1 by reconfiguring two photodiode service mezzanines, MezzPD0 and MezzPD1.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×